OpenStudy (anonymous):

Find the absolute maximum and minimum values of the following function on the given interval. If there are multiple points in a single category list the points in increasing order in x value and enter N in any blank that you don't need to use. f(x)=9x^3−27x+9, [0,3]

OpenStudy (anonymous):

i know that you have to find the directive that is 27^2-27

OpenStudy (anonymous):

I keep getting the wrong answer (1,0) and (-1,0) the right answer is (3,171) and (1,-9) I just don't know how to get that to work out

OpenStudy (zehanz):

The derivative (not directive ;)) is\[f'(x)=27x^2-27\]\[27x^2-27=0 \Leftrightarrow x^2-1=0 \Leftrightarrow x^2=1 \Leftrightarrow x= \pm 1\] You have also found these x-values, but you plugged them in f', which will give you 0. You should put these x-values in the original function.

OpenStudy (anonymous):

that gave me -9 and 27

OpenStudy (zehanz):

\[f(-1)=9(-1)^3-27 \cdot -1 + 9=-9+27+9=27\]\[f(1)=9 \cdot 1^3 -27 \cdot 1+9=9-27+9=-9\]Rembember, these outcomes, 27 and -9 are -at this moment- only the y-coordinates of points of the graph of f where there is a horizontal tangent line (f'= 0). We do not know if they are minimum of maximum values. Because the domain is [0, 3], the graph begins in x=0 and ends at x=3. You also have to know what f(0) and f(3) are.

OpenStudy (anonymous):

I got 9 and 171

OpenStudy (zehanz):

f(0)=9 \[f(3)=9 \cdot 3^3 -27 \cdot 3 + 9=171\] This is the graph: So: x = -1 is irrelevant: ouside of [0,3]. Absolute minimum for x = 1, value -9 Absolute maximum for x = 3, value 171

OpenStudy (anonymous):

where did you get the number 1

OpenStudy (zehanz):

Although f(0)=9, 9 has no special meaning. it is just an in-between value. They ask you for the absolute minumum and the abs max. Because graphs go up and down, the points with a horizontal tangent line are interesting: with this function, you got a horizontal tangent line for x = 1. If you look at the graph, you'll see that for x=1, the graph is at its lowest point (-9). The highest point is not where the tangent line is horizontal, it's just where the quickly ascending graph is cut off because it goes no further than x=3. Value there: 171.

OpenStudy (zehanz):

You'll have to realize that this graph is only part of the complete graph: it's the part that goesfrom x=0 to x=3. See image. You can also see the two values, -1 and , where the tangent line is horizontal. For x=-1 there is a (local) maximum, for x=1 there is a (local) minimum.